OpenStudy (anonymous):

Hey !! I need help with 4 questions please. If you can help me I will fan and medal! :)

OpenStudy (anonymous):

https://gyazo.com/0f4737ab4597cfc4665b76ca28ebca7c

pooja195 (pooja195):

start by adding 5 to both sides

OpenStudy (anonymous):

-10 and 18

pooja195 (pooja195):

when u add 5 to -5 u get 0 so they cancel out! :) 18 is right now we have x/6=18 Now multiply by the reciprocal (flip x/6 to 6/x)

pooja195 (pooja195):

18 x 6 = ?

OpenStudy (anonymous):

108

pooja195 (pooja195):

Good so x=108

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

So that's the answer

pooja195 (pooja195):

yep

OpenStudy (anonymous):

Thanks :)

pooja195 (pooja195):

:)

OpenStudy (anonymous):

Are you helping me with the other 3?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

ye

OpenStudy (anonymous):

Ok https://gyazo.com/53c9973fcb6bace0b1dacaf570e61b6d

pooja195 (pooja195):

Start by multiplying 3 from both sides

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

12 and 24

pooja195 (pooja195):

\[\huge~\rm~\frac{ (12+4n)\times 3 }{ 3} =8 \times 3 \]

pooja195 (pooja195):

ye so 12+4n=24 now solve for x :)

OpenStudy (anonymous):

384

pooja195 (pooja195):

No... 24-12=?

OpenStudy (anonymous):

I'm sorry

OpenStudy (anonymous):

12

pooja195 (pooja195):

12/4=value of x

OpenStudy (anonymous):

3

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

Is that right

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I will post another question so you can get your medal again ok? @pooja195

pooja195 (pooja195):

I dont care about medals i care about helping :)

OpenStudy (anonymous):

Okay something happened to my computer btw